OpenStudy (anonymous):

Solve: (1/3)(6 – 3t) + 18 = 0

OpenStudy (anonymous):

this is in the wrong section.. you should go into the math switch that were all the math People are:)

OpenStudy (anonymous):

(2-t) + 18=0 20-t=0 20= -t -20= t